As the director of capital budgeting for KU dairy, you are evaluating two mutually exclusive projects...

  1. As the director of capital budgeting for KU dairy, you are evaluating two mutually exclusive projects with the following net cash flows:

Year

Project A Cash Flow

Project B Cash Flow

0

-100,000

-100,000

1

50,000

10,000

2

40,000

30,000

3

30,000

40,000

4

10,000

60,000

If KU’s cost of capital is 15%, examine which project you would choose?
Answer #1

We will choose the project which has the higher NPV

Project A:

NPV = Present value of cash inflows - present value of cash outflows

NPV = -100,000 + 50,000 / (1 + 0.15)1 + 40,000 / (1 + 0.15)2 + 30,000 / (1 + 0.15)3 + 10,000 / (1 + 0.15)4

NPV = -100,000 + 43,478.26087 + 30,245.74669 + 19,725.48697 + 5,717.53246

NPV = -832.97

Project B:

NPV = Present value of cash inflows - present value of cash outflows

NPV = -100,000 + 10,000 / (1 + 0.15)1 + 30,000 / (1 + 0.15)2 + 40,000 / (1 + 0.15)3 + 60,000 / (1 + 0.15)4

NPV = -100,000 + 8,695.6522 + 22,684.31002+ 26,300.6493 + 34,305.19474

NPV = -8,014.19

Both projects should be REJECTED as both have negative NPVs